Conner is 30 years younger than Noelle. 6 years ago, Noelle's age was 3 times Conner's age. How old is Conner now?
Answer by JulietG(1812)

C = N-30
N-6 = 3(C-6)
Distribute
N-6 = 3C-18
Add 6 to each side
N = 3C - 12
Substitute the value of C from the top equation into the bottom
N = 3(N-30)-12
Distribute
N = 3N - 90 - 12
N = 3N - 102
Add 102 to each side
N + 102 = 3N
Subtract N from each side
102 = 2N
Divide each side by 2
51 = N
If Noelle is 51, then Conner is 30 less than that, or 21

Prove: 6 years ago, Noelle's age was 3 times Conner's age
6 years ago, Noelle was 51-6, or 45.
6 years ago, Conner was 21-6, or 15.
45 is 3x 15.
Conner is currently 21, and Noelle is currently 51.